The mean birth weight for Ethiopian babies is μ =3.1 kg with sd (σ=0.5 kg)
A) What weight values encompass the 95% newborns?
B) If the cut of point for low birth weight is 2.5kg, what proportion of the new
borns are lbw?
C) Find the 10th percentile birth weight.

Answer by Boreal(15235)

z=(x-mean)/sd
95% are between +/- 1.96 z or 0.98 kg
so the weight is 3.1 +/- 0.98 or (2.12 kg, 4.08 kg)
-
z(2.5)=(2.5-3.1)/0.5
=-1.2
Probability z < -1.2 is 0.1151
-
z(0.10)=-1.282
-1.282=(x-3.1)/0.5
x=2.46 kg